A rug is made by weaving 20 lb of yarn that is 50% wool with a yarn that is 25% wool. How many pounds of the yarn that is 25% wool must be used if the finished rug is to be 35% wool?
Answer by richwmiller(17219)

We have 20 lb of 50% wool
We want to add some 25% wool to make 35% wool
0.25*x+0.5*20=0.35(20+x)
0.25*x+10=7+0.35x
0.25x-0.35x=7-10
-0.1x=-3
x=30 lb at 25% (7.5) is added to 20 lb at 50%(10) to make 35% of 50 lb(17.5)
check
0.25*30+0.5*20=0.35(20+30)
7.5+10=0.35(50)
17.5=17.5
ok